A Bibliography of Journals Accepting Manuscripts Written using TEX

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu (Internet)
WWW URL: https://www.math.utah.edu/~beebe/

30 November 2023 Version 1.55

Abstract

This bibliography lists journals that accept papers written using TEX, and use TEX to produce the output page images without rekeying the text.

Title word cross-reference

Academy [DOK, IZV, SBO]. Acids [NAR]. Algebra [J-La]. Algebraic [JAL]. American [BAM, JAMb, MEM, NOT, PAM, TAM]. Applications [J-La]. Applied [APP, JAMa, QAM]. Approximate [INT]. Atmospheric [ACP02].

B [NPB, PLB, PRB]. Biology [J-C94, JMB, JTB]. Bulletin [BAM, CMB].

C [JCL, PRC]. Canadian [CJM, CMB]. Carolinae [COM]. Chemistry [ACP02]. Commentationes [COM]. Communications [CPC]. Computation [APP, JSY85, MATa]. Computational [J-C94, THE]. Computations [INT91]. Computer [JCo92, TCS, CGF, CPC]. Current [CMP].

D [PRD, P-D]. Design [EPO ]. Differential [JDI]. Dissemination [EPO ]. Doklady [DOK]. dynamics [THE].

Electronic [EPO , Sol92]. Energy [JHE97]. Expositions [SUG].

fluid [THE]. fluids [PREb]. Forum [CGF].

Geography [Sol92]. Geometry [JAL, JDI]. Glaciology [JG]. Graphics [CGF].

High [JHE97].

Imprint [PI]. Informatica [INFb]. Information [INFa]. Institute [PRO]. interdisciplinary [PREb]. International [INT]. Interval [INT91]. Izvestiya [IZV].

Journal [CJM, INT, J-C94, J-Lb, JCo92, JHE97, NYJ94, Sol92, JAMa, JP, JCL, JG, JMB, JAL, JAMb, JDI, JSY85, JTB, MJ, STP].

Language [JCL]. Letters [PLA, PLB]. Linear [J-La]. Logic [J-Lb].

Maple [MAP]. Mathematica [MJ]. Mathematicae [COM]. Mathematical [BAM, CMB, CMP, JAMb, MEM, NOT, PAM, STP, TOP, TAM, TRM, MATb]. Mathematics [APP, NYJ94, Sol92, CJM, DOK, IZV, MATa, PRO, QAM, SBO]. matter [PREa]. Mechanics [JAMa]. Memoirs [MEM]. Molecular [JMB]. Moscow [TRM].

Newsletter [MAP]. nonlinear [PREa]. Notices [NOT]. Nuclear [NPa, NPB]. Nucleic [NAR].

Origination [EPO ].

Petersburg [STP]. Philosopher [PI]. Philosophica [RES13]. Philosophy [RP]. Physica [P-D]. Physical [PRA, PRB, PRC, PRD, PREb, PREa]. Physics [CPC, JHE97, NPa, NPB, PREb, ACP02, JP, PLA, PLB, PRec, PREa]. Physik [ZP]. plasmas [PREb]. Probability [TOP]. Proceedings [PAM, PRO]. Programming [J-Lb]. Publications [CMP]. Publishing [EPO ].

Quarterly [QAM].

Radical [RP]. Reasoning [INT]. related [PREb]. Reports [PRec]. Res [RES13]. Research [NAR]. Review [PRA, PRB, PRC, PRD, PREb, PREa]. Reviews [MATb]. Russian [DOK, IZV, SBO].

Sbornik [SBO]. Science [TCS]. Sciences [DOK, INFa, IZV, SBO]. Security [JCo92]. Society [BAM, JAMb, MEM, NOT, PAM, TAM, TRM]. soft [PREa]. Solstice [Sol92]. St. [STP]. Statistical [PREb, PREa]. Statistics [TOP]. Steklov [PRO]. Sugaku [SUG]. Symbolic [JSY85].

Technical [MAP]. TEXline [TEX ]. Theoretical [JTB, THE, TCS]. Theory [TOP]. topics [PREb]. Transactions [TAM, TRM]. Translation [JCL].

Universitatis [COM].

York [NYJ94].

Zeitschrift [ZP].

References

tj-acp

[ACP02] Atmospheric Chemistry and Physics, 2002. ISSN 1680-7316. URL http:/

                                                /www.atmos-chem-phys.org/technical-

tex.htm. European Geophysical Society. Typeset with pdfL^{A}TEX. Authors are strongly encouraged to submit manuscripts in L^{A}TEX using the egs.cls style file. As a reward, page charges will be lower than for submission in other formats. In the first step, a pdf file will be produced suitable for on-line viewing. After an Internet-based discussion of the scientific contents the paper will (if accepted) be formatted in the traditional two-column layout using the same L^{A}TEX source.

tj-appl-math-comp

[APP] Applied Mathematics and Computation. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). North-Holland Publishing Company, Amsterdam, The Netherlands. From volume 59, number 1, November 1993, the journal cover proclaims in a fancy logo “Now accepting L^{A}TEX manuscripts”. Contact Victor van Beuren, Editor, Physical Sciences Journals Group, Elsevier Science Inc., 655 Avenue of the Americas, New York, NY 10010, e-mail: vvb@panix.com, for further details.

tj-bull-ams

[BAM] Bulletin of the American Mathematical Society. CODEN BAMOAD. ISSN 0273-0979 (print), 1088-9485 (electronic). American Mathematical Society, Providence, RI, USA. Use of AMS-TEX began in 1983; completely in AMS-TEX and AMS-L^{A}TEX since late 1986; style files

available at e-math.ams.org.

tj-cgf

[CGF] Computer Graphics Forum. CO-DEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic). North-Holland Publishing Company, Amsterdam, The Netherlands. Publication switched from troff to TEX in summer 1992.

tj-can-j-math

[CJM] Canadian Journal of Mathematics. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic). Completely done in TEX since about 1988, including covers and index. The Bulletin consists of four issues of 144 pages each per year, while the Journal consists of six issues of 224 pages each. The final output is produced from PostScript files, so the particular variation of TEX (L^{A}TEX, AMS-TEX, plain) isn’t important, as long as the source document can be adapted to conform to the style of each journal. A CMSTEX style file for plain TEX is used internally, and a L^{A}TEX style file is nearly complete.

tj-can-math-bull

[CMB] Canadian Mathematical Bulletin. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic). The Canadian Mathematical Society employs TEX for their typesetting, and encourages authors to submit that way. See also the note under [CJM].

tj-cmp

[CMP] Current Mathematical Publications. CODEN CUMPBW. ISSN 0361-4794. American Mathematical Society, Providence, RI, USA. Prepared completely with TEX since 1985; issue indexes prepared with TEX from 1983.

tj-comment-math

[COM] Commentationes Mathematicae Universitatis Carolinae. CODEN CMU-CAA. ISSN 0010-2628. Charles University, Praha, Czechoslovakia. Published with AMS-TEX since 1990.

Articles using amsppt.sty are accepted.

tj-cpc

[CPC] Computer Physics Communications.

CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). Elsevier Science Publishers, Amsterdam, The Netherlands. L^{A}TEX format accepted for special proceedings issues.

tj-dokl-math

[DOK] Russian Academy of Sciences. Doklady. Mathematics. ISSN 1064-5624 (print), 1531-8362 (electronic). American Mathematical Society, Providence, RI, USA. Translation into English of the entire pure mathematics section of the “Doklady Rossiĭskoĭ Akademiĭ Nauk”. (Formerly “Soviet Mathematics— Doklady”.) Prepared with AMS-TEX beginning in 1983.

tj-epodd

[EPO ] David F. Brailsford and Richard K.

Furuta, editors. Electronic Publishing—Origination, Dissemination, and Design, 1988–. CODEN EPODEU. ISSN 0894-3982. Wiley, New York, NY, USA. The publisher provides style files for TEX, L^{A}TEX, and troff article formats.

tj-info-sci

[INFa] Information Sciences. CODEN ISI-JBC. ISSN 0020-0255 (print), 1872-6291 (electronic). North-Holland Publishing Company, Amsterdam, The Netherlands. Contact Victor van Beuren, Editor, Physical Sciences Journals Group, Elsevier Science Inc., 655 Avenue of the Americas, New York, NY 10010, e-mail: vvb@panix.com, for further details.

tj-informatica

[INFb] Informatica. CODEN INFOFF. ISSN 0350-5596. The Slovene Society Informatika, Ljubljana, Slovenia. Typeset with L^{A}TEX. Authors can submit articles using the informat style option, which is printed in its entirety in the back of each issue.

tj-int-j-approx-reasoning

[INT] International Journal of Approximate Reasoning. CODEN IJARE4. ISSN 0888-613X. North-Holland Publishing Company, Amsterdam, The Netherlands. Official publication of the North American Fuzzy Information Processing Society. Vol. 1, no. 1 (January 1987) –. Contact Victor van Beuren, Editor, Physical Sciences Journals Group, Elsevier Science Inc., 655 Avenue of the Americas, New York, NY 10010, email: vvb@panix.com, for further details.

tj-intcom

[INT91] Vyacheslav M. Nesterov, editor. Interval Computations, 1991. ISSN 0135-4868. Box 52, St.Petersburg 195256, Russia, E-mail: nest@nit. spb.su (Internet). A refereed international journal, Interval Computations is the only periodical in the world devoted specifically to various aspects of reliable numerical computations based on the interval approach. It is managed by an international editorial board from Bulgaria, Germany, Japan, Russia, and the United States, and printed in Russia. The journal includes various items in the fields of theoretical research, computer tools, applications, interdisciplinary research and other relevant areas. The IC has been prepared with TEX since 1992. Recently a new style has been designed based on standard L^{A}TEX article style with changes in fonts, headings layout, running heads etc. It uses 10.8pt Baskerville for body text (the unusual font size is selected for better wedding with CM math) and horizontally-squeezed Palatino for headings. For mathematical formulae, standard Computer Modern is preserved. Proofs are made using emTEX DVIHPLJ on a Hewlett– Packard LaserJet IIIp, and final pages are printed at 600dpi. Using VTEX is now under consideration. Papers are accepted in any TEX variety; L^{A}TEX article.sty preferred.

tj-izv-math

[IZV] Russian Academy of Sciences.

Izvestiya. Mathematics. American Mathematical Society, Providence, RI, USA. Translation into English of “Izvestiya Rossiĭskoĭ Akademiĭ Nauk Seriya Matematicheskaya”. (Formerly “Mathematics of the USSR— Izvestiya”.) Prepared with AMS-TEX since 1986.

tj-j-comput-biol

[J-C94] Journal of Computational Biology, 1994. CODEN JCOBEM. ISSN 1066-5277. URL http://

www.catchword.com/rpsv/cw/mal/
10665277/contp1.htm. Mary Ann

Liebert Inc., 2 Madison Avenue, Larchmont, NY 10538-1962, USA.

TEX is the preferred form for manuscript submission.

tj-j-lin-alg-appl

[J-La] Linear Algebra and Its Applications. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). North-Holland Publishing Company, Amsterdam, The Netherlands. Vol. 1 (January 1968) –. Contact Victor van Beuren, Editor, Physical Sciences Journals Group, Elsevier Science Inc., 655 Avenue of the Americas, New York, NY 10010, e-mail: vvb@panix.com, for further details.

tj-j-logic-prog

[J-Lb] The Journal of Logic Programming. CODEN JLPRE2. ISSN 0743-1066. North-Holland Publishing Company, Amsterdam, The Netherlands. Vol. 1, no. 1 (June 1984) –. Contact Victor van Beuren, Editor, Physical Sciences Journals Group, Elsevier Science Inc., 655 Avenue of the Americas, New York, NY 10010, e-mail: vvb@panix.com, for further details.

tj-jnl-alg-geom

[JAL] Journal of Algebraic Geometry. ISSN 1056-3911. University Press, Inc., Naperville, IL, USA.

tj-j-app-mech

[JAMa] Journal of Applied Mechanics. CO-DEN JAMCAV. ISSN 0021-8936. TEX macros available from Charles Beardsley, Managing Director of Publications, ASME, Tel: +1 212 705-7722.

tj-jnl-ams

[JAMb] Journal of the American Mathematical Society. ISSN 0894-0347. American Mathematical Society, Providence, RI, USA. Prepared with AMS-TEX and AMS-L^{A}TEX; style files available at e-math.ams.org.

tj-jclt

[JCL] Rex Jaeschke, editor. Journal of C Language Translation. ISSN 1042-5721. Vol. 1, no. 1 (June 1989)–. Ceased with vol. 3, no. 4, published in 1992.

tj-jcs

[JCo92] Journal of Computer Security, 1992.

IOS Press, Amsterdam, The Netherlands. Style files are available from the journal editors.

tj-jnl-diff-geom

[JDI] Journal of Differential Geometry.

CODEN JDGEAS. ISSN 0022-040X. Lehigh University, Bethlehem, PA, USA. Prepared with TEX since 1988.

tj-jg

[JG] Journal of Glaciology. CODEN JOGLAO. ISSN 0022-1430. British Glaciological Society, Cambridge, UK. First issue typeset by TEX in 1991.

tj-jhep

[JHE97] Journal of High Energy Physics, 1997. CODEN JHEPFG, JHEPAB (paper archive). ISSN 1029-8479.

URL http://jhep.mse.jhu.edu/.

Societa italiana di fisica, Bologna, Italy. Online journal. The paper version of the journal was discontinued at the end of 2000, but publication continues in electronic form. The master archive is hosted at the “Florence National Library, which is the Italian body officially in charge of keeping the legal archive of all editorial products published in Italy, both electronic and paper, thereby ensuring their legal va-

lidity.” See http://jhep.mse.jhu.
edu/JOURNAL/IoPP_SISSA.html for

details. Articles are marked up exclusively in TEX, with PostScript for figures, and are available in DVI, PDF, and PostScript formats.

tj-jmb

[JMB] Journal of Molecular Biology. CO-DEN JMOBAK. ISSN 0022-2836. Academic Press, New York, NY, USA. L^{A}TEX and BibTEX style files

jmb.sty and jmb.bst have been

provided by Tom Schneider (e-mail

toms@ncifcrf.gov).

tj-j-phys

[JP] Journal of Physics. Plain TEX macros available.

tj-jsc

[JSY85] Bruno Buchberger, editor. Journal of Symbolic Computation, 1985. ISSN 0747-7171 (print), 1095-855X (electronic). Academic Press, New York, NY, USA. A L^{A}TEX style file for the journal is available from the publisher, and from ftp.risc.

uni-linz.ac.at in the directory /
pub/jsc in the files jsc.sty and

jscguide.tex. The resolution is 600 dpi. The software used is PCTEX.

tj-jtb

[JTB] Journal of Theoretical Biology. CO-DEN JTBIAP. ISSN 0022-5193. L^{A}TEX and BibTEX style files jmb. sty and jmb.bst have been provided by Tom Schneider (e-mail toms@

ncifcrf.gov).

MAPLETECH

[MAP] Tony Scott, editor. The Maple Technical Newsletter. ISSN 1061-5733. Birkhäuser, Cambridge, MA, USA; Berlin, Germany; Basel, Switzerland. Published twice annually.

tj-math-comp

[MATa] Mathematics of Computation. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic). American Mathematical Society, Providence, RI, USA. Use of AMS-TEX began in 1983; completely in AMS-TEX and AMS-L^{A}TEX since 1988; style files available at e-math.

ams.org.

tj-math-rev

[MATb] Mathematical Reviews. CODEN MAREAR. ISSN 0025-5629. American Mathematical Society, Providence, RI, USA. Beginning in November 1979, TEX has been used for issue indexes; since January 1985, the entire journal has been prepared in TEX. The data for this journal is archived with several database services and is also available on CD-ROM as MathSci.

tj-memo-ams

[MEM] Memoirs of the American Mathematical Society. CODEN MAM-CAU. ISSN 0065-9266. American Mathematical Society, Providence, RI, USA. Copy is author prepared; AMS-TEX and AMS-L^{A}TEX style files available at e-math.ams.

org.

tj-mathematica-j

[MJ] Stephen Wolfram, editor. Mathematica Journal. ISSN 1047-5974 (print), 1097-1610 (electronic). Advanced Book Program, Addison-Wesley Publishing Company, 350 Bridge Parkway, Redwood City, CA 94065. Published both electronically and in print using TEX.

tj-nar

[NAR] Nucleic Acids Research. CODEN NARHAD. ISSN 0305-1048. L^{A}TEX and BibTEX style files nar.sty and nar.bst have been provided by Tom Schneider (e-mail: toms@ncifcrf.

                                        gov).

tj-noti-ams

[NOT] Notices of the American Mathematical Society. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic). American Mathematical Society, Providence, RI, USA. Use of TEX began in 1982 with selected sections; completely in TEX from 1988 to 1994. Since 1995, the journal has been typeset by Quark Express, although author manuscripts in TeX are accepted, and automatically converted to Quark format at the publisher.

tj-np-a

[NPa] Nuclear Physics A. CODEN NU-PABL. ISSN 0375-9474. Elsevier Science Publishers, Amsterdam, The Netherlands. L^{A}TEX format accepted.

tj-np-b

[NPB] Nuclear Physics B. CODEN NUPBBO. ISSN 0550-3213 (print), 1873-1562 (electronic). Elsevier Science Publishers, Amsterdam, The Netherlands. L^{A}TEX format accepted.

tj-nyjm

[NYJ94] Mark Steinberger, editor. The New York Journal of Mathematics, 1994. ISSN 1076-9803. University at Albany, State University of New York, Albany, NY, USA. Electronic submission and publication only. The New York Journal of Mathematics is a refereed mathematics journal being launched by the University at Albany, State University of New York.

The journal is broadly based in subject matter, covering algebra, modern analysis, and geometry/topology.

The papers will be presented via the TEX mathematical typesetting language, developed by Donald Knuth. TEX has become the standard for computer typesetting of mathematical publications.

Access will be provided through a variety of electronic means, including listserv, anonymous ftp and gopher. In particular, access will be much more convenient than that provided by a conventional print journal. Because papers will be published as they are accepted by the editorial board, publication delays will be minimal.

The articles will be archived, date stamped, and maintained by the University at Albany, State University of New York. The University takes responsibility for insuring the integrity of the journal’s archive in perpetuity.

The journal will have an ISSN number and will be accessed by the major bibliographic publications in the field.

The development of this journal is being underwritten by the Office of Information Systems of the University at Albany, State University of New York.

The New York Journal will be accessible by a combination of listserv

                                        and gopher/ftp.

By listserv, we will distribute abstracts of the papers, together with instructions for retrieving them via

                                        ftp and gopher.

The abstracts will be available on four separate listserv lists: one will distribute abstracts for every paper. The other three are specialty lists, giving abstracts for papers in algebra, analysis and geometry/ topology, respectively. The reader may, of course, subscribe to any combination of these lists.

tj-physica-d

[P-D] Physica D. CODEN PDNPDT. ISSN 0167-2789 (print), 1872-8022 (electronic). Elsevier Science Publishers, Amsterdam, The Netherlands. L^{A}TEX format accepted for special proceedings issues.

tj-proc-ams

[PAM] Proceedings of the American Mathematical Society. CODEN PAMYAR. ISSN 0002-9939 (print), 1088-6826 (electronic). American Mathematical Society, Providence, RI, USA. Use of AMS-TEX began in 1983; completely in AMS-TEX and AMS-L^{A}TEX since late 1986; style files

available at e-math.ams.org.

tj-pi

[PI] Philosopher’s Imprint. ISSN 1533-

628X.

tj-phys-letters-a

[PLA] Physics Letters A. CODEN PY-LAAG. ISSN 0375-9601 (print), 1873-2429 (electronic). Elsevier Science Publishers, Amsterdam, The Netherlands. L^{A}TEX format accepted.

tj-phys-letters-b

[PLB] Physics Letters A and B. CODEN PYLBAJ. ISSN 0370-2693. Elsevier Science Publishers, Amsterdam, The Netherlands. L^{A}TEX format accepted.

tj-phys-rev-a

[PRA] Physical Review A. CODEN PLRAAN. ISSN 1050-2947 (print), 1094-1622, 1538-4446, 1538-4519. American Physical Society, Ridge, NY 11961, USA. L^{A}TEX format accepted.

tj-phys-rev-b

[PRB] Physical Review B. CODEN PRB-MDO. ISSN 0163-1829. American Physical Society, Ridge, NY 11961, USA. L^{A}TEX format accepted.

tj-phys-rev-c

[PRC] Physical Review C. CODEN PRV-CAN, PRDOF8. ISSN 0556-2813, 1089-4918. American Physical Society, Ridge, NY 11961, USA. L^{A}TEX format accepted.

tj-phys-rev-d

[PRD] Physical Review D. CODEN PRV-DAQ. ISSN 0556-2821 (print), 1089-4918 (electronic), 1538-4500. American Physical Society, Ridge, NY 11961, USA. L^{A}TEX format accepted.

tj-phys-rev-e2

[PREa] Physical Review E, Statistical, nonlinear, and soft matter physics. CODEN PRESCM. ISSN 1539-3755 (print), 1550-2376 (electronic).

URL http://pre.aps.org. Amer-

ican Physical Society, Ridge, NY 11961, USA. L^{A}TEX format accepted.

tj-phys-rev-e1

[PREb] Physical Review E, Statistical physics, plasmas, fluids, and related interdisciplinary topics. CODEN PLEEE8. ISSN 1539-3755 (print), 1550-2376 (electronic). American Physical Society, Ridge, NY 11961, USA. L^{A}TEX format accepted.

tj-phys-repd

[PRec] Physics Reports. CODEN PRPLCM. ISSN 0370-1573. Elsevier Science Publishers, Amsterdam, The Netherlands. L^{A}TEX format accepted on an experimental basis.

tj-proc-steklo

[PRO] Proceedings of the Steklov Institute of Mathematics. ISSN 0081-5438. American Mathematical Society, Providence, RI, USA. Translation into English of the “Trudy ordena Lenina Matematicheskogo instuta imeni V. A. Steklova”. Prepared with AMS-TEX began since 1987.

tj-quart-app-math

[QAM] Quarterly of Applied Mathematics. CODEN QAMAAY. ISSN 0033-569X. Brown University, Providence, RI, USA.

tj-res

[RES13] Res Philosophica, 2013. ISSN 2168-9105 (print), 2168-9113 (electronic).

tj-rp

[RP] Radical Philosophy. ISSN 0300-211X.

tj-sbor-math

[SBO] Russian Academy of Sciences.

Sbornik. Mathematics. ISSN 1064-5616. American Mathematical Society, Providence, RI, USA. Translation into English of “Matematicheskiĭ Sbornik”. (Formerly “Mathematics of the USSR—Sbornik”.) Prepared with AMS-TEX since 1987.

tj-solstice

[Sol92] Solstice: An Electronic Journal of Geography and Mathematics, 1992. ISBN 1-877751-44-8; 1-877751-52-9; 1-877751-53-7. ISSN 1059-5325. Institute of Mathematical Geography, 2790 Briarcliff, Ann Arbor, MI 48105-1429, USA.

tj-stp-math-j

[STP] St. Petersburg Mathematical Journal. ISSN 1061-0022. American Mathematical Society, Providence, RI, USA. Translation into English of “Algebra i Analiz”. (Formerly “Leningrad Mathematical Journal”.) Prepared with AMS-TEX.

tj-sugaku

[SUG] Sugaku Expositions. ISSN 0898-9583. American Mathematical Society, Providence, RI, USA. Translation into English of expository articles from “Sugaku”. Prepared with AMS-TEX.

tj-trans-ams

[TAM] Transactions of the American Mathematical Society. CODEN TAM-TAM. ISSN 0002-9947 (print), 1088-6850 (electronic). American Mathematical Society, Providence, RI, USA. Use of AMS-TEX began in 1983; author submission of AMS-TEX manuscripts encouraged since 1987; AMS-TEX and AMS-L^{A}TEX since late 1986; style files available

at e-math.ams.org.

tj-tcs

[TCS] Theoretical Computer Science. CO-DEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

URL http://www.elsevier.com:
80/inca/publications/store/5/

0/5/6/. Elsevier Science Publishers, Amsterdam, The Netherlands. L^{A}TEX format accepted.

tj-texline

[TEX ] Malcolm Clark, editor. TEXline, 1987–. ISSN 0961-3978. This is an informal newsletter of the TEX community.

tj-theor-comput-fluid-dyn

[THE] Y. M. Hussaini, editor. Theoretical and computational fluid dynamics. CODEN TCFDEP. ISSN 0935-4964. URL http:

//link.springer-ny.com/link/
service/journals/00162/index.

htm. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc. Volume 1, number 1, appeared in 1989, and authors are encouraged to submit articles in L^{A}TEX form. Tables of contents are available online for issues beginning with volume 8, number 1, 1996.

tj-thy-prob-stat

[TOP] Theory of Probability and Mathematical Statistics. CODEN TPM-SCO. ISSN 0094-9000 (print), 1547-7363 (electronic). American Mathematical Society, Providence, RI, USA. Translation into English of “Teoria Veroyatnostei i Matematicheskaya Statistika”. Prepared with AMS-TEX.

tj-trans-moscow

[TRM] Transactions of the Moscow Mathematical Society. CODEN TMMSD4. ISSN 0077-1554. American Mathematical Society, Providence, RI, USA. Translation into English of “Trudy Moskovskogo Matematicheskogo Obshchestva”. Prepared with AMS-TEX since 1987.

tj-z-phys

[ZP] Zeitschrift für Physik. Need to confirm this.